# Does Worst-Performing Agent Lead the Pack? Analyzing Agent Dynamics in Unified Distributed SGD

### v1

@@ 2 Preliminaries
where $\Delta_{n}\triangleq\sum_{s=0}^{n-1}(\mathbf{g}(X_{s})-\mathbb{E}_{\boldsymbol{\pi}}(\mathbf{g}))$. By following the algebraic manipulations in [13, Theorem 6.3.7] for asymptotic variance (univariate version), we can rewrite (4) in a matrix form such that
@@ 3 Asymptotic Analysis of UD-SGD
Theorem 3.2 is achieved by decomposing the Markovian noise term $\nabla F_{i}(\theta_{n}^{i},X_{n}^{i})-\nabla f_{i}(\theta_{n}^{i})$, using the Poisson equation technique as discussed in [7, 30, 18], into a Martingale difference noise term, along with additional noise terms. We then reformulate (6) into an iteration akin to stochastic approximation, as depicted in (56). The subsequent step involves verifying the conditions on these noise terms under our stated assumptions. Crucially, this theorem also establishes that UD-SGD ensures an almost sure convergence of each agent to an optimal point $\theta^{*}\in\mathcal{L}$, even in scenarios where the communication interval $K_{l}$ gradually increases, in accordance with Assumption 2.3-ii). The detailed proof of this theorem is provided in Appendix C.
Central Limit Theorem: Let $\mathbf{U}_{i}\!\triangleq\!\boldsymbol{\Sigma}_{X^{i}}(\nabla F_{i}(\theta^{*},\cdot))$ represent the asymptotic covariance matrix (defined in (5)) associated with each agent $i\in[N]$, given their sampling strategy $\{X^{i}_{n}\}$ and function $\nabla F_{i}(\theta^{*},\cdot)$. Define $\mathbf{U}\triangleq\frac{1}{N^{2}}\sum_{i=1}^{N}\mathbf{U}_{i}$. We assume the polynomial step-size $\gamma_{n}\!\sim\!\gamma_{\star}/n^{a}$, $a\!\in\!(0.5,1]$ and $\gamma_{\star}>0$. In the case of $a=1$, we further assume $\gamma_{\star}>1/2\mu$, where $\mu$ is defined in (8).
@@
The proof, presented in Appendix D, addresses technical challenges in deriving the CLT for UD-SGD, specifically the second-order conditions in decomposing the Markovian noise term, which is not present in the i.i.d. sampling case [58, 43, 51]. This decomposition needs to be carefully designed to separate the consensus error $\nabla F_{i}(\theta_{n}^{i},X_{n}^{i})\!-\!\nabla F_{i}(\theta_{n},X_{n}^{i})$, which is analyzed in terms of Lemma 3.1, from the bias $\nabla F_{i}(\theta_{n},X_{n}^{i})\!-\!\nabla f_{i}(\theta_{n})$ out of the Markovian sampling. The specifics of this analysis are expanded upon in Appendices D.1 to D.3. We require $\gamma_{\star}\!>\!1/2\mu$ for $a\!=\!1$ to ensure that the largest eigenvalue of $\mathbf{M}$ is negative, as this is a necessary condition for the existence of $\mathbf{V}$ in (12) (otherwise integration diverges). In the case where there is only one agent ($N\!=\!1$), $\mathbf{V}$ and $\mathbf{V}^{\prime}$ reduce to the matrices specified in the CLT result of vanilla SGD [30, 38, 52]. In addition, for a special case of constant communication interval in Assumption 2.3-i) and i.i.d. sampling as shown in Table 1, we recover the CLT of LSGD-RC in [51]. See Appendix E for detailed discussions.
@@ 4 Experiments
where the feature vector $\mathbf{x}_{i,j}$ and its corresponding label $y_{i,j}$ are held by agent $i$, with a penalty parameter $\kappa$ set to $1$. We use the ijcnn1 dataset [15] with $22$ features in each data point and $50$k data points in total, which is evenly distributed to two groups with $50$ agents each ($N=100$ agents in total) and each agent holds $B=500$ data points. Each agent in the first group has full access to its entire dataset, and thus can employ i.i.d. sampling (baseline) or single shuffling. On the other hand, each agent in the other group has a graph-like structure and uses SRW (baseline), NBRW or SRRW with reweighting to sample its local dataset with uniform weight. In this simulation, we assume that agents can only communicate through a communication network using the DSGD algorithm. This scenario with heterogeneous agents, as depicted in Figure 1, is of great interest in large-scale machine learning [37, 33]. In addition, we employ a decreasing step size $\gamma_{n}=1/n$ in our UD-SGD framework (1) because it is typically used for the strongly convex objective function and is tested to have the fastest convergence in this simulation setup. Due to space constraints, we defer detailed simulation setup, including the introduction of SRW, NBRW, and SRRW, to Section G.1.
@@
Figure 2: Binary classification problem. From left to right: (a) Impact of efficient sampling strategies on convergence. (b) Performance gains from partial adoption of efficient sampling. (c) Comparative advantage of SRRW over NBRW in a small subset of agents.
@@
The simulation results are obtained through $120$ independent trials. In Figure 2, we assume that the first group of agents perform shuffling method while the other group of agents all change their sampling strategies from baseline SRW to NBRW and SRRW, as shown in the legend. This plot shows that improved sampling strategy leads to overall convergence speedup since NBRW and SRRW are more efficient than SRW [38, 26]. Furthermore, it illustrates that SRRW is significantly more efficient than NBRW in this simulation setup, i.e., SRRW $\gg$ NBRW $>$ SRW in terms of sampling efficiency. The comparison between i.i.d. sampling and the shuffling method in the first group of agents (while keeping the second group of agents unchanged) is demonstrated in Figure 3 in Appendix G.1, where shuffling method outperforms i.i.d. sampling. In Figure 2, we let the first group of agents perform i.i.d. sampling while only changing a portion of agents in the second group to upgrade from SRW to SRRW, e.g., $30$ SRW $20$ SRRW in the legend means that there are 30 agents using SRW while the rest $20$ agents in the second group upgrade to SRRW. We observe that more agents willing to upgrade from SRW to SRRW lead to smaller asymptotic MSE, as predicted from Theorem 3.3 and Remark 3. This improvement in MSE reduction doesn’t scale linearly with more agents adopting SRRW because each agent holds its own dataset that are not necessarily identical, resulting in different individual limiting covariance matrices $\mathbf{V}_{i}\!\neq\!\mathbf{V}_{j}$. While maintaining i.i.d. sampling for the first group of agents, we compare the performance when the second group of agents in Figure 2 employ NBRW or SRRW. Remarkably, the case with only $10$ agents out of $50$ agents in the second group adopting far more efficient sampling strategy ($40$ SRW, $10$ SRRW) through incentives or compliance already produces a smaller MSE than all $50$ agents using slightly better strategy ($50$ NBRW). The performance gap becomes even more pronounced when $20$ agents upgrade from SRW to SRRW ($30$ SRW, $20$ SRRW). We show that the performance of a distributed system can be improved significantly when a small proportion of agents adopt highly efficient sampling strategies.
@@
Furthermore, in Section G.2, we simulate an image classification task with CIFAR-10 dataset [44] by training a $5$-layer neural network collaboratively through a $10$-agent network. The result is illustrated in Figure 4, where SRRW outperforms NBRW and SRW as anticipated. In summary, we find that upgrading even a small portion of agents to efficient sampling strategies (e.g., shuffling method, NBRW, SRRW under different dataset structures) improves system performance in UD-SGD. These results are consistent in binary and image classification tasks, underscoring that every agent matters in distributed learning.
@@ 5 Conclusion
In this work, we develop an UD-SGD framework that establishes the CLT of various distributed algorithms with Markovian sampling. We overcome the technical challenges such as quantifying consensus error under very general communication patterns and decomposing Markovian noise through the Poisson equation, which extends the analysis beyond the single-agent scenario. We demonstrate that even if only a few agents optimize their sampling strategies, the entire distributed system will benefit with smaller limiting covariance in the CLT, suggesting a reduced MSE. This finding challenges the current established upper bounds where the worst-performing agent leads the pack. Future studies could pivot towards developing fine-grained finite-time bounds to individually characterize each agent’s behavior, and theoretically analyze the effect of SRRW in UD-SGD.
@@ F.2 Discussion on partial client sampling
As far as we are aware, the FedAvg algorithm under the Markovian sampling strategy remains unexplored in the FL literature. Extrapolating the behavior of unselected agents in FedAvg from i.i.d sampling to Markovian sampling suggests that unselected agents would remain idle. In contrast, our algorithm enables unselected agents to continue evolving $X^{i}_{n}$. These additional transitions contribute to faster mixing of the Markov chain for each unselected agent and a smaller bias of $F_{i}(\theta,X^{i}_{n})$ relative to its mean-field $f_{i}(\theta)$, potentially accelerating the convergence.
@@ G.1 Simulation Setup in Section 4
This simulation is performed on a PC with an R9 5950x CPU, RTX 3080 GPU and 128 GB RAM. In this simulation, we assume that agents follow the DSGD algorithm (1). Specifically, all agents are distributed over a communication network. In order to create this network among $100$ agents and the graph-like dataset structure held by each agent, we utilize connected sub-graphs from the real-world graph Facebook in SNAP [49]. All $100$ agents collaborate together to generate a deterministic communication matrix $\mathbf{W}=[W_{ij}]$ with Metropolis Hasting algorithm of the following form: For $i,j\in[N]$, we have
@@
Figure 3: Comparison of various sampling strategies in binary classification problem. The 1-sigma error bar is the standard deviation of the MSE $\mathbb{E}[\|\theta_{n}-\theta^{*}\|^{2}]$ from $120$ independent experiments, calculated through numpy.std() method.
@@
    where matrix $\mathbf{P}=[P_{ij}]$ is the transition kernel of the baseline Markov chain and $\boldsymbol{\mu}=[\mu_{i}]$ is its corresponding stationary distribution. Additionally, $\alpha$ denotes the force of self repellence, and larger $\alpha$ leads to stronger force of self repellence, thus higher sampling efficiency [26, Corollary 4.3]. Moreover, vector $\mathbf{x}\in\mathbb{R}^{N}$ is in the interior of probability simplex, representing the empirical distribution, in other words, the visit frequency of each node in the graph. The update rule of this empirical distribution is in the following form:
@@
Now, we perform the simulation in which all agents in each group are free to change their sampling strategies, e.g., first group of $50$ agents can pick either i.i.d. sampling or shuffling method and second group of $50$ agents can select either SRW, NBRW, or SSRW, thus resulting in six scenarios in total. As illustrated in Figure 3, when the second group of agents fix their sampling strategies, we can see that shuffling method outperforms i.i.d. sampling with smaller asymptotic MSE. However, shuffling method may not perform perfectly for small time $n$ due to slow mixing behavior in the initial period, which is also observed in the single-agent scenario in [65, 2, 38]. The error bar therein also indiates that the random-walk sampling strategy has a significant impact on the overall system performance and SRRW has smaller variance than NBRW and SRW.
@@ G.2 Image Classification Task
Figure 4: Comparison of various sampling strategies in image classification problem using $5$-layer neural network.

### v2

@@ 2 Preliminaries
where $\Delta_{n}\triangleq\sum_{s=0}^{n-1}(\mathbf{g}(X_{s})-\mathbb{E}_{\boldsymbol{\pi}}(\mathbf{g}))$. By following the algebraic manipulations in [12, Theorem 6.3.7] for asymptotic variance (univariate version), we can rewrite (4) in a matrix form such that
@@ 3 Asymptotic Analysis of UD-SGD
Theorem 3.2 is achieved by decomposing the Markovian noise term $\nabla F_{i}(\theta_{n}^{i},X_{n}^{i})-\nabla f_{i}(\theta_{n}^{i})$, using the Poisson equation technique as discussed in [6, 29, 17], into a Martingale difference noise term, along with additional noise terms. We then reformulate (6) into an iteration akin to stochastic approximation, as depicted in (56). The subsequent step involves verifying the conditions on these noise terms under our stated assumptions. Crucially, this theorem also establishes that UD-SGD ensures an almost sure convergence of each agent to a local minimum $\theta^{*}\in\mathcal{L}$, even in scenarios where the communication interval $K_{l}$ gradually increases, in accordance with Assumption 2.3-ii). The detailed proof of this theorem is provided in Appendix C.
Central Limit Theorem: Let $\mathbf{U}_{i}\!\triangleq\!\boldsymbol{\Sigma}_{X^{i}}(\nabla F_{i}(\theta^{*},\cdot))$ represent the asymptotic covariance matrix (defined in (5)) associated with each agent $i\in[N]$, given their sampling strategy $\{X^{i}_{n}\}$ and function $\nabla F_{i}(\theta^{*},\cdot)$. Define $\mathbf{U}\triangleq\frac{1}{N^{2}}\sum_{i=1}^{N}\mathbf{U}_{i}$. We assume the polynomial step-size $\gamma_{n}\!\sim\!\gamma_{\star}/n^{a}$, $a\!\in\!(0.5,1]$ and $\gamma_{\star}>0$. In the case of $a=1$, we further assume $\gamma_{\star}>1/2\mu$, where $\mu$ is defined in (8). For notational simplicity, and without loss of generality, our remaining CLT result is stated while conditioning on the event that $\{\theta_{n}\to\theta^{*}\}$ for some $\theta^{*}\in\mathcal{L}$.
@@
The proof, presented in Appendix D, addresses the technical challenges in deriving the CLT for UD-SGD, specifically the second-order conditions in decomposing the Markovian noise term, which is not present in the i.i.d. sampling case [58, 43, 51]. We decompose $\nabla F_{i}(\theta_{n},X_{n}^{i})\!-\!\nabla f_{i}(\theta_{n})$ into three parts in (48) using Poisson equation: $e_{n+1}^{i},\nu_{n+1}^{i},\xi_{n+1}^{i}$. The consensus error $\theta_{n}^{i}-\theta_{n}$ embedded in noise terms $e_{n+1}^{i}$ and $\xi_{n+1}^{i}$ is a new factor, whose characteristics have been quantified in our Lemma 3.1 but are not present in the single-agent scenario analyzed as an application of stochastic approximation in [22, 29]. The specifics of this analysis are expanded upon in Appendices D.1 to D.3. We require $\gamma_{\star}\!>\!1/2\mu$ for $a\!=\!1$ to ensure that the largest eigenvalue of $\mathbf{M}$ is negative, as this is a necessary condition for the existence of $\mathbf{V}$ in (12) (otherwise integration diverges). In the case where there is only one agent ($N\!=\!1$), $\mathbf{V}$ and $\mathbf{V}^{\prime}$ reduce to the matrices specified in the CLT result of vanilla SGD [29, 38, 52]. In addition, for a special case of constant communication interval in Assumption 2.3-i) and i.i.d. sampling as shown in Table 1, we recover the CLT of LSGD-RC in [51]. See Appendix E for detailed discussions.
@@ 4 Experiments
where the feature vector $\mathbf{x}_{i,j}$ and its corresponding label $y_{i,j}$ are held by agent $i$, with a penalty parameter $\kappa$ set to $1$. We use the ijcnn1 dataset [14] with $22$ features in each data point and $50$k data points in total, which is evenly distributed to two groups with $50$ agents each ($N=100$ agents in total) and each agent holds $B=500$ distinct data points. Each agent in the first group has full access to its entire dataset, and thus can employ i.i.d. sampling (baseline) or single shuffling. On the other hand, each agent in the other group has a graph-like structure and uses SRW (baseline), NBRW or SRRW with reweighting to sample its local dataset with uniform weight. In this simulation, we assume that agents can only communicate through a communication network using the DSGD algorithm. This scenario with heterogeneous agents, as depicted in Figure 1, is of great interest in large-scale machine learning [37, 32]. In addition, we employ a decreasing step size $\gamma_{n}=1/n$ in our UD-SGD framework (1) because it is typically used for the strongly convex objective function and is tested to have the fastest convergence in this simulation setup. Due to space constraints, we defer detailed simulation setup, including the introduction of SRW, NBRW, and SRRW, to Section G.1.
@@
Figure 2: Binary classification problem. From left to right: (a) Impact of efficient sampling strategies on convergence. (b) Performance gains from partial adoption of efficient sampling. (c) Comparative advantage of SRRW over NBRW in a small subset of agents. (d) Asymptotic network independence of four algorithms under UD-SGD framework with fixed sampling strategy (shuffling, SRRW). (e) Different sampling strategies in the DSGD algorithm with time-varying topology (DSGD-VT). (f) Different sampling strategies in the DFL algorithm with increasing communication interval.
@@
The simulation results are obtained through $120$ independent trials. In Figure 2, we assume that the first group of agents perform either i.i.d. sampling or shuffling method, while the other group of agents all change their sampling strategies from baseline SRW to NBRW and SRRW, as shown in the legend. This plot shows that improved sampling strategy leads to overall convergence speedup since NBRW and SRRW are more efficient than SRW [38, 25]. Furthermore, it illustrates that SRRW is significantly more efficient than NBRW in this simulation setup, i.e., SRRW $\gg$ NBRW $>$ SRW in terms of sampling efficiency. While keeping the second group of agents unchanged, we can see that shuffling method outperforms i.i.d. sampling with smaller asymptotic MSE. However, shuffling method may not perform perfectly for small time $n$ due to slow mixing behavior in the initial period, which is also observed in the single-agent scenario in [65, 1, 38]. The error bar therein also indicates that the random-walk sampling strategy has a significant impact on the overall system performance and SRRW has smaller variance than NBRW and SRW.
In Figure 2, we let the first group of agents perform i.i.d. sampling while only changing a portion of agents in the second group to upgrade from SRW to SRRW, e.g., $30$ SRW $20$ SRRW in the legend means that there are 30 agents using SRW while the rest $20$ agents in the second group upgrade to SRRW. We observe that more agents willing to upgrade from SRW to SRRW lead to smaller asymptotic MSE, as predicted by Theorem 3.3 and Remark 3. This improvement in MSE reduction doesn’t scale linearly with more agents adopting SRRW because each agent holds its own dataset that are not necessarily identical, resulting in different individual limiting covariance matrices $\mathbf{V}_{i}\!\neq\!\mathbf{V}_{j}$.
While maintaining i.i.d. sampling for the first group of agents, we compare the performance when the second group of agents in Figure 2 employ NBRW or SRRW. Remarkably, the case with only $10$ agents out of $50$ agents in the second group adopting far more efficient sampling strategy ($40$ SRW, $10$ SRRW) through incentives or compliance already produces a smaller MSE than all $50$ agents using slightly better strategy ($50$ NBRW). The performance gap becomes even more pronounced when $20$ agents upgrade from SRW to SRRW ($30$ SRW, $20$ SRRW). We show that the performance of a distributed system can be improved significantly when a small proportion of agents adopt highly efficient sampling strategies.
Figure 2 empirically illustrates the asymptotic network independence property via four algorithms under our UD-SGD framework: Centralized SGD (communication interval $K=1$, communication matrix $\mathbf{W}=\mathbf{1}\mathbf{1}^{T}/N$); LSGD-FP (FL with full client participation, $K=5$, $\mathbf{W}=\mathbf{1}\mathbf{1}^{T}/N$); DSGD-VT (DSGD with time-varying topologies, randomly chosen from $5$ doubly stochastic matrices); DFL (decentralized FL with fixed MH-generated $\mathbf{W}$ and increasing communication interval $K_{l}=\max\{1,\log(l)\}$ after $l$-th aggregation). We fix the sampling strategy (shuffling, SRRW) throughout this plot. All four algorithms overlap around $1000$ steps, implying that they have entered the asymptotic regime with similar performance where the CLT result dominates, implying the asymptotic network independence in the long run.
Figure 2 and 2 show the performance of different sampling strategies in DSGD-VT and DFL algorithms in terms of MSE. Both plots consistently demonstrate that improving agent’s sampling strategies (e.g., shuffling > iid sampling, and SRRW > NBRW > SRW) leads to faster convergence with smaller MSE, supporting our theory.
@@
Furthermore, in Section G.2, we simulate an image classification task with CIFAR-10 dataset [44] by training a $5$-layer CNN and ResNet-18 model collaboratively through a $10$-agent network. The result is illustrated in Figure 3, where SRRW outperforms NBRW and SRW as expected. In summary, we find that upgrading even a small portion of agents to efficient sampling strategies (e.g., shuffling method, NBRW, SRRW under different dataset structures) improves system performance in UD-SGD. These results are consistent in binary and image classification tasks, underscoring that every agent matters in distributed learning.
@@ 5 Conclusion
In this work, we develop an UD-SGD framework that establishes the CLT of various distributed algorithms with Markovian sampling. We overcome technical challenges such as quantifying consensus error under very general communication patterns and decomposing Markovian noise through the Poisson equation, which extends the analysis beyond the single-agent scenario. We demonstrate that even if only a few agents optimize their sampling strategies, the entire distributed system will benefit with a smaller limiting covariance in the CLT, suggesting a reduced MSE. This finding challenges the current established upper bounds where the worst-performing agent leads the pack. Future studies could pivot towards developing fine-grained finite-time bounds to individually characterize each agent’s behavior, and theoretically analyze the effect of SRRW in UD-SGD.
6 Acknowledgments and Disclosure of Funding
We thank the anonymous reviewers for their constructive comments. This work was supported in part by National Science Foundation under Grant Nos. CNS-2007423, IIS-1910749, and IIS-2421484.
@@ F.2 Discussion on partial client sampling
To our knowledge, the FedAvg algorithm under the Markovian sampling strategy remains unexplored in the FL literature. Extrapolating the behavior of unselected agents in FedAvg from i.i.d sampling to Markovian sampling suggests that unselected agents would remain idle. In contrast, our algorithm enables unselected agents to continue evolving $X^{i}_{n}$. These additional transitions contribute to faster mixing of the Markov chain for each unselected agent and a smaller bias of $F_{i}(\theta,X^{i}_{n})$ relative to its mean field $f_{i}(\theta)$, potentially accelerating the convergence.
@@ G.1 Simulation Setup in Section 4
This simulation is performed on a PC with an AMD R9 5950X, RTX 3080 and 128 GB RAM. In this simulation, we assume that agents follow the DSGD algorithm (1). In Figure 2 - 2, each agent holds a disjoint local dataset (non-overlapping data points for every agent), while we distribute the ijcnn1 dataset [14] with more varied distribution among $100$ agents by leveraging Dirichlet distribution with the default alpha value of $0.5$ in Figure 2 - 2.
Moreover, we assume that all agents are distributed over a communication network. In order to create this network among $100$ agents and the graph-like dataset structure held by each agent, we utilize connected sub-graphs from the real-world graph Facebook in SNAP [49]. All $100$ agents collaborate together to generate a deterministic communication matrix $\mathbf{W}=[W_{ij}]$ with Metropolis Hasting algorithm of the following form: For $i,j\in[N]$, we have
@@
    where matrix $\mathbf{P}=[P_{ij}]$ is the transition kernel of the baseline Markov chain and $\boldsymbol{\mu}=[\mu_{i}]$ is its corresponding stationary distribution. Additionally, $\alpha$ denotes the force of self repellence, and larger $\alpha$ leads to stronger force of self repellence, thus higher sampling efficiency [25, Corollary 4.3]. Moreover, vector $\mathbf{x}\in\mathbb{R}^{N}$ is in the interior of probability simplex, representing the empirical distribution, in other words, the visit frequency of each node in the graph. The update rule of this empirical distribution is in the following form:
@@ G.2 Image Classification Task
Figure 3: Image classification experiment. From left to right: (a) Comparison of various sampling strategies in image classification problem using $5$-layer neural network. (b) Train a $5$-layer CNN model with different number of total agents (clients) to show the linear speedup effect. (c) Train ResNet-$18$ model with different sampling strategies among $10$ agents with participation ratio $0.4$.
@@
In Figure 3 and 3, we perform image classification experiments in the FL setting with partial client participation. Only $4$ random agents will participate in the training process at each aggregation phase. In Figure 3, we fix the sampling strategy (shuffling, SRRW with $\alpha=10$) and test the linear speedup effect for the $5$-layer CNN model by duplicating $10$ agents to $N$ agents with $N\in\{10,20,30,40\}$, keeping the same participation ratio $0.4$. As can be seen from the plot, the training loss is inversely proportional to the number of agents, i.e., at $200$ rounds, the training loss is $0.52$ for $10$ agents, $0.23$ for $20$ agents, $0.18$ for $30$ agents, and $0.12$ for $40$ agents. In Figure 3, we extend the current simulation from 5-layer CNN model to ResNet-18 model [34] in order to numerically test the performance of different sampling strategies in a more complex neural network training task. By fixing the shuffling in the first group of agents, we observe that improving Markovian sampling from SRW to NBRW, then to SRRW, gives accelerated training process with smaller training loss.